suppose that 1 out of 10 plasma televisions shipped with a defective speaker. out of a shipment of n=400 plasma televisions. find the probability (as a %) that there are more than 28 with defective speakers
1
Expert's answer
2020-04-30T17:40:59-0400

Let "X=" the number of plasma televisions with defective speakers: "X\sim Bin(n, p)"

Given "n=400, p=0.1"

Use the dishonest-coin principle with p=0.1 to find


"\mu=np=400(0.1)=40"

"\sigma=\sqrt{np(1-p)}=\sqrt{400(0.1)(1-0.1)}=6"

"28=40-2\times6=\mu-2\sigma"

By the 68–95–99.7 rule


"Pr(\mu-2\sigma\leq X\leq \mu+2\sigma)\approx 0.9545"

The probability that there are more than 28 with defective speakers


"0.9545+{1-0.9545\over 2}\approx0.97725"

Or

"n=400\geq30, np=40\geq5, n(1-p)=360\geq5"

Then, "X"  has an approximately normal distribution with mean "\mu=np=40" and standard deviation 

"\sigma=\sqrt{np(1-p)}=6"


"P(X>28)=P(X>28+0.5)=1-P(X\leq28.5)="

"1-P(Z\leq{28.5-40\over6}) \approx1-0.02764\approx0.97236"

By Binomial distribution


"P(X>28)\approx0.976515"